# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ B.2 Proofs of theoretical results for the Classical error model
Since $\operatorname{\mathbb{P}}_{Y}^{n}$ does not depend on $X$, it admits the trivial conditional mean embedding operator $\mathcal{C}_{Y}:\operatorname{\mathcal{H}_{k_{X}}}\rightarrow\operatorname{\mathcal{H}_{k_{Y}}}$ defined by $\mathcal{C}_{Y}f=\mu_{\operatorname{\mathbb{P}}^{n}_{Y}}$ for any $f\in\operatorname{\mathcal{H}_{k_{X}}}$. Hence, for any $f\in\operatorname{\mathcal{H}_{k_{X}}}$, and using Assumption 4.0.2, we have:
$\sup_{y_{1:n}}\|\mu_{\operatorname{\mathbb{P}}^{n}_{Y}}\|^{2}_{\operatorname{\mathcal{H}_{k_{Y}}}}=\sup_{y_{1:n}}\left\|\frac{1}{n}\sum_{i=1}^{n}k_{Y}(y_{i},\cdot)\right\|^{2}_{\operatorname{\mathcal{H}_{k_{Y}}}}=\sup_{y_{1:n}}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}k_{Y}(y_{i},y_{i})+\frac{2}{n^{2}}\sum_{i\neq j}k_{Y}(y_{i},y_{j})\right)=1$
hence $\sup_{y_{1:n}}\|\mu_{\operatorname{\mathbb{P}}^{n}_{Y}}\|_{\operatorname{\mathcal{H}_{k_{Y}}}}\leq 1$ and $\|\mathcal{C}_{Y}\|_{o}\leq 1$. Therefore,
$\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\mathbb{P}^{0,n},\mathbb{P}_{XY})]]$ $\leq\|\mathcal{C}_{Y}\|_{o}\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{X})]]$
$\leq\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{n},\operatorname{\mathbb{P}}_{X})]]$
$\leq\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X})+\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X}^{n})]]$
$\leq\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X})]]+\frac{1}{\sqrt{n}}$
$\leq\frac{c+2}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})$
$\quad\quad+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})$
where we used Lemma B.3 for the first inequality, the triangle inequality for the third and Lemma B.6 with $k_{X}^{2}$ for the last inequality. The fact that $k_{X}^{2}$ satisfies the required conditions of Lemma B.6 follows from Assumptions 4.0.2 and 4.0.3. In particular, if $|k_{X}|\leq 1$ then $|k^{2}_{X}|\leq 1$ and if $k_{X}$ is translation invariant then $k^{2}_{X}(x,x^{\prime})=\psi(x-x^{\prime})\psi(x-x^{\prime})\forall x,x^{\prime}\in\mathcal{X}$ and $k^{2}_{X}$ is also translation invariant. Finally, for the fourth inequality, we used Lemma B.1 with $S_{i}=(x_{i})\sim Q_{i}:=\operatorname{\mathbb{P}}^{0}_{X}$. ∎
Lemma B.8.
@@
(Classical) Let $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}_{X}^{0}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ and $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$. Under Assumptions 1-4, for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0}\operatorname{\mathbb{P}}_{g(\theta,\cdot)},\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)})]]$
@@
$\quad\quad\quad\leq\Lambda\left(\frac{1}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right)$
@@
Using Lemma B.3 and Assumption 4 we have that for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}),\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}]]$ $\leq\Lambda\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X})]]$
and the result follows by Lemma B.6 for $k^{2}_{X}$. ∎
@@
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})$ $\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
@@
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
@@
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta}^{0},\operatorname{\mathbb{P}}_{\theta})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
@@
where we used the triangle inequality in the first two and last two inequalities and the definition of $\theta^{*}(\operatorname{\mathbb{P}})$ in the third inequality. Using lemma B.1 with $S_{i}=(x_{i},y_{i})\sim Q_{i}:=\operatorname{\mathbb{P}}^{0}$ we have that:
@@
Using equation (47) and lemmas B.7 and B.8, we have that:
@@
$\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\frac{1}{\sqrt{n}(c+1)}+\sqrt{\frac{1}{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})\right.$
$\left.\quad\quad\quad\quad+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right).$
B.2.2 Proof of Corollary 4.2.2 (Classical)

### v3

@@ D.2 Proofs of theoretical results for the Classical error model
The proof follows exactly the proof of the Berkson case in Lemma D.4 with a few changes to account for the different error structure. To avoid repetition, we only note the differences. The only part that depends on the ME structure is the bound on the term
$\sum_{i=1}^{n}\sqrt{\operatorname{\mathbb{E}}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mu_{\operatorname{\mathbb{F}}_{X\mid w_{i}}}\|^{2}_{\mathcal{H}_{k_{X}}}]}$
as this depends directly on the structure of the prior centring measure $\operatorname{\mathbb{F}}_{w_{i}}$ relative to the true distribution of each $(x_{i},w_{i})$. For every $i\in\{1,\dots,n\}$ we have:
$\operatorname{\mathbb{E}}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mu_{\operatorname{\mathbb{F}}_{X\mid w_{i}}}\|^{2}_{\mathcal{H}_{k_{X}}}]$
$=\operatorname{\mathbb{E}}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mathbb{E}_{x\sim\operatorname{\mathbb{F}}_{X\mid w_{i}}}[\phi_{X}(x)]\|^{2}]$
$=\operatorname{\mathbb{E}}_{x_{i}\sim\operatorname{\mathbb{P}}^{0}}[k(x_{i},x_{i})]-2\operatorname{\mathbb{E}}_{x_{i}\sim\operatorname{\mathbb{P}}^{0}}[\left<\phi_{X}(x_{i}),\operatorname{\mathbb{E}}_{x\sim\operatorname{\mathbb{F}}_{X\mid w_{i}}}\phi_{X}(x)\right$>]
$\quad\quad+\left<\operatorname{\mathbb{E}}_{x\sim\operatorname{\mathbb{F}}_{w_{i}}}[\phi_{X}(x)],\operatorname{\mathbb{E}}_{x\sim\operatorname{\mathbb{F}}_{X\mid w_{i}}}[\phi_{X}(x)]\right>_{\mathcal{H}_{k_{X}}}$
$=\operatorname{\mathbb{E}}_{x_{i}\sim\operatorname{\mathbb{P}}^{0}}[k(x_{i},x_{i})]-\left<\mu_{\operatorname{\mathbb{P}}_{X}^{0}},\mu_{\operatorname{\mathbb{P}}_{X}^{0}}\right>_{\mathcal{H}_{k_{X}}}+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}_{X\mid w_{i}})$
$=\text{Var}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{P}}_{X}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}_{X\mid w_{i}})$
and hence
$\sum_{i=1}^{n}\sqrt{\operatorname{\mathbb{E}}_{w_{i},x_{i}}[\|\phi_{X}(x_{i})-\mu_{\operatorname{\mathbb{F}}_{X\mid w_{i}}}\|^{2}_{\mathcal{H}_{k_{X}}}]}$
$\quad=\sum_{i=1}^{n}\sqrt{\text{Var}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{P}}_{X}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}_{X\mid w_{i}})}.$
Unsurprisingly, this term now depends on the RKHS variance of the true marginal of $X$ as well as the quality of the prior centring measure with respect to $\operatorname{\mathbb{P}}_{X}^{0}$. This is because in the Classical error case, $X$ is not dependent on the ME, and hence if $X$ has a large variance, then $W$ will have a larger variance and the problem becomes harder. The rest of the proof is identical to Lemma D.4. ∎
Lemma D.8.
@@
(Classical) Let $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}_{X}^{0}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ and $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ and suppose Assumptions 4.0.1-4.0.4 hold, then for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)},\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)})]]$
@@
$\quad\leq\Lambda\left(\frac{1}{\sqrt{n}(c+1)}+\frac{1}{\sqrt{n(c+2)}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right)$
@@
Using Lemma D.3 and Assumption 4 we have that for any $\theta\in\Theta$:
$\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{X}^{0}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}),\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}]]$
$\quad\leq\Lambda\mathbb{E}_{x_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{P}}_{X})]]$
and the result follows by Lemma D.6 for $k^{2}_{X}$. ∎
@@
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
@@
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{XY},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
$\quad\quad+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
@@
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta}^{0},\operatorname{\mathbb{P}}_{\theta})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
@@
where we used the triangle inequality in the first three and last two inequalities and the definition of $\theta^{\star}(\operatorname{\mathbb{P}})$ in the third inequality. Since this holds for any $\theta\in\Theta$ we have that
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})$ $\leq\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0}_{\theta})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta}^{0},\operatorname{\mathbb{P}}_{\theta})$
$\quad+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
$\quad+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY}).$
Using lemma D.1 with $S_{i}=(x_{i},y_{i})\sim Q_{i}:=\operatorname{\mathbb{P}}^{0}_{X\>|\>w_{i}}\times\operatorname{\mathbb{P}}_{Y\>|\>x}^{0}$ we have that:
@@
Using equation (41) and lemmas D.7 and D.8, we have that:
@@
$\quad\quad\leq\frac{1}{\sqrt{n}}\left(\frac{2\Lambda}{c+1}+\frac{2\Lambda+2}{\sqrt{c+2}}\right)$
$\quad\quad+\frac{2c}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{n}\sum_{i=1}^{n}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{P}}_{X}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}_{w_{i}})}\right)$
$\quad\quad+\frac{2}{c+1}\left(\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\sqrt{2}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}\right).$
D.2.2 Proof of Corollary 4.2.2 (Classical)